# Fusion Systems related to polynomial representations of $\operatorname{SL}_{2}(q)$

### v1

@@ 7. Uniqueness of the saturated fusion systems on $S$
Lastly suppose that $S\cong S_{\Lambda}(q)$ and ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$. Let $k_{R}$ be the lift to $\mathrm{Aut}_{\mathcal{F}}(S)$ of an element of order $q-1$ in $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))$ so that $k_{R}$ acts on $V$. Let $k$ to be any element of order $q-1$ in $N_{\mathrm{Aut}_{\mathcal{F}}(V)}(\mathrm{Aut}_{S}(V))$ in the same Hall $p^{\prime}$-subgroup of $N_{\mathrm{Aut}_{\mathcal{F}}(V)}(\mathrm{Aut}_{S}(V))$ as $k_{R}|_{V}$, whose lift $k_{V}$ to $\mathrm{Aut}_{\mathcal{F}}(S)$ normalises $R=S_{0}Z(S)$. Then $K_{R}:=\langle k_{R}\rangle$ normalises $K_{V}:=\langle k_{V}\rangle$ and so to prove that $|\mathrm{Out}_{\mathcal{F}}^{0}(S)|=(q-1)^{2}$ as in (4), it suffices to verify $K_{R}\cap K_{V}=1$. By Proposition 6.3, $V$ is isomorphic to $\Lambda(q)$ as an $\mathbb{F}_{p}O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$-module. Let $W$ be the unique proper non-trivial $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$-submodule of $V$. Then $\mathrm{Aut}_{\mathcal{F}}(S)$ normalises $W$ since $V$ is characteristic in $S$ and $\mathrm{Aut}_{\mathcal{F}}(S)|_{V}$ normalises $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$. Let $t\in K_{R}\cap K_{V}$. Then since $t$ is the lift of an element of $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))$, $t$ acts trivially on $C_{V}(S)W/W=C_{R}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R)))W/W\cong C_{R}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R)))$. On the other hand, since $V/W$ is isomorphic as a $\mathbb{F}_{p}O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$-module to a natural $\operatorname{SL}_{2}(q)$-module, every non-trivial $p^{\prime}$-element in $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$ acts fixed point freely on $V/W$. This implies that $t=1$ so that $K_{R}\cap K_{V}=1$, as desired.
@@
Let $\mathfrak{D}$ be a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}(S)$ containing a Hall $p^{\prime}$-subgroup $K$ of $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$. Conjugating by $\mathrm{Aut}(S)$ if necessary we may assume that $\mathrm{Aut}_{\mathcal{F}}^{0}(S)\leq\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}(S)$ and that $\mathfrak{D}$ is a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}(S)$. Recall that the map $\delta:\mathrm{Aut}(S)\to\Delta=(\mathrm{Aut}(S,S/V),\mathrm{Aut}(S,Z(S)))$ of Lemma 2.20 is injective upon restriction to $\mathfrak{D}$. Thus in each case, there is $k\in\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ cyclic of order $q-1$ determined by the lift of an element in $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(X))}(\mathrm{Aut}_{S}(X))$, for $X\in\{Q,R\}$ which determines a unique subgroup of $\mathfrak{D}$. Similarly if $V\in\mathcal{E}(\mathcal{F})$, then the image under $\delta$ of lift to $S$ of an element of $\frac{q-1}{(n,2)}$ in $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))}(\mathrm{Aut}_{S}(V))$ uniquely determines the actions on $S/V$ and $Z(S)$. We conclude that $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ is uniquely determined. ∎
@@
If $S=S_{n}(q)$ for some $12\leqslant n\leqslant p-$ 1 and $\mathcal{E}(\mathcal{F})=R^{S}$ then $\mathcal{F}$ is isomorphic to a $p^{\prime}$-index subsystem of $\mathcal{F}^{*}(n,q,R)_{P}$.
@@
In all remaining cases, we may assume that $V\in\mathcal{E}(\mathcal{F})$ by Propositions 6.1 and 6.2.
@@
It thus remains to treat the case when $p=3$ and $S=S_{\Lambda}(q)$. Observe that $\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V)\cong\mathrm{GL}_{2}(q)$ and so $O^{p}(N_{\mathrm{Aut}(V)}(\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V)))\leq\mathrm{Aut}_{\mathcal{F}_{\Lambda}^{*}(q)}(V)C_{\mathrm{Aut}(V)}(\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V))$. it thus suffices to show that $C:=C_{\mathrm{Aut}(V)}(\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V))=Z(\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V))$ has order $q-1$. Denote by $W$ the unique proper non-trivial $\mathbb{F}_{3}\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V)$-submodule of $V$. Let $s$ be a $3$-element in $C$. Then $[V,s]$ and $[W,s]$ are both normalised by $\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V)$ and since $[V,s]W<V$ and $[W,s]<W$, we must have $[V,s]\leq W$ and $[W,s]=1$. Let $K$ be a Hall $3^{\prime}$-subgroup of $N_{\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V)}(\mathrm{Aut}_{S}(V))$ and define $K_{1}=C_{K}([V,S]/W)$, a group of order $q-1$. The Three Subgroups Lemma implies that $[s,[V,S],K_{1}]=1$ so that $[s,[V,S]]\leq C_{W}(K_{1})=1$, where the last equality follows from Lemma 2.16. Hence, $[V,S]\leq C_{V}(s)$. But $C_{V}(s)$ is clearly normalised by $\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V)$, and we deduce that $V=C_{V}(s)$ and $s=1$. Hence, $C$ is a $3^{\prime}$-group. But $Z(\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V))\leq C$ and $Z(\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V))$ acts transitively on non-trivial vectors in $C_{V/W}(S)$. Using Lemma 5.16 we conclude that $C=C_{C}(C_{V/W}(S))Z(\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V))$. An application of the A×$\times$×B-Lemma [Asc86, (24.2)] gives that $C_{C}(C_{V/W}(S))$ centralises $V/W$. But then $C_{C}(C_{V/W}(S))$ centralises $\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V)$ and so $C_{V}(C_{C}(C_{V/W}(S)))$ is a $\mathbb{F}_{3}\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V)$-submodule of $V$. Since $V$ is indecomposable, $C_{C}(C_{V/W}(S))$ centralises $V$ so $C_{C}(C_{V/W}(S))=1$, and $C=Z(\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(V))$ has order $q-1$, as desired. ∎
@@
If $S=S_{n}(q)$ and ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup Q^{S}$ then $O^{p^{\prime}}(\mathcal{F})\cong O^{p^{\prime}}(\mathcal{F}^{*}(n,q,Q))$ and $\mathcal{F}$ is isomorphic to a $p^{\prime}$-index subsystem of $\mathcal{F}^{*}(n,q,Q)$.
@@
Write $X:=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(Q))$ and $X^{*}:=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,Q)}(Q))$ so that $N_{X}(\mathrm{Aut}_{S}(Q))=N_{X^{*}}(\mathrm{Aut}_{S}(Q))$. By Proposition 2.5 and Corollary 2.9, $\mathrm{Aut}(Q)=CH$ where $H\cong\Gamma\mathrm{L}_{2}(q)$ and $C$ is a normal $p$-subgroup of central automorphisms of $Q$. It follows immediately that $XC=X^{*}C$. Let $\tau$ be an involution in $N_{X}(\mathrm{Aut}_{S}(Q))$, so $X=\mathrm{Inn}(Q)C_{X}(\tau)$ and $X^{*}=\mathrm{Inn}(Q)C_{X^{*}}(\tau)$. Since $[C_{C}(\tau),Q,\tau]=1$, the Three Subgroups Lemma implies that $[Q,C_{C}(\tau)]=[[\tau,Q]Z(Q),C_{C}(\tau)]=1$, and so $C_{C}(\tau)=1$. We deduce that $C_{X}(\tau)=C_{XC}(\tau)=C_{X^{*}C}(\tau)=C_{X^{*}}(\tau)$, so $X=X^{*}$ and $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(Q))=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,Q)}(Q))$, as desired. ∎
@@
If $S=S_{n}(q)$ and ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$ then $O^{p^{\prime}}(\mathcal{F})\cong O^{p^{\prime}}(\mathcal{F}^{*}(n,q,R))$ and $\mathcal{F}$ is isomorphic to a $p^{\prime}$-index subsystem of $\mathcal{F}^{*}(n,q,R)$.
@@
By Propositions 7.5 and 7.6 it suffices to show that $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ uniquely determines $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))$. Set $X:=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))$ and $X^{*}=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}(R))$ and observe that $N_{X}(\mathrm{Aut}_{S}(R))$ is contained in the restriction of $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ to $R$. Upon demonstrating that $N_{X}(\mathrm{Aut}_{S}(R))=N_{X^{*}}(\mathrm{Aut}_{S}(R))$, the result will follow from Lemma 5.17. Recall that $\mathrm{Aut}_{\mathcal{F}}^{0}(S)=\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}^{0}(S)$ and let $k$ be the lift of an element of order $q-1$ in $N_{X}(\mathrm{Aut}_{S}(R))$ to $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$. Then $k$ acts as $\mu^{-1}$ on $S/V$ and as $\mu$ on $C_{V}(S)$. By Lemma 2.20, $\delta$ is injective upon restricting to a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}^{0}(S)$, so the restriction of $k$ to $R$ lies in $N_{X^{*}}(\mathrm{Aut}_{S}(R))$. But then $N_{X}(\mathrm{Aut}_{S}(R))=\langle\mathrm{Aut}_{S}(R),k\rangle=N_{X^{*}}(\mathrm{Aut}_{S}(R))$ and so Lemma 5.17 applies. This completes the proof. ∎
@@
If $S=S_{\Lambda}(q)$ and ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$ then $O^{p^{\prime}}(\mathcal{F})\cong O^{p^{\prime}}(\mathcal{F}^{*}_{\Lambda}(q))$ and $\mathcal{F}$ is isomorphic to a $p^{\prime}$-index subsystem of $\mathcal{F}^{*}_{\Lambda}(q)$.
@@
By Propositions 7.5 and 7.6 it suffices to show that $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ uniquely determines $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))$. Let $k$ be the lift to $\mathrm{Aut}_{\mathcal{F}}(S)$ of an element of $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))}(\mathrm{Aut}_{S}(R))$ of order $q-1$. Since $\mathrm{Aut}_{\mathcal{F}}^{0}(S)=\mathrm{Aut}_{\mathcal{F}^{*}_{\Lambda}(q)}^{0}(S)$, $k$ lies in a Hall $p^{\prime}$-subgroup of $N_{\mathrm{Aut}_{\mathcal{F}^{*}_{\Lambda}(q)}^{0}(S)}(R)$ and there is a subgroup $K_{V}$ of order $q-1$ satisfying
@@
Thus $K_{V}$ restricts faithfully to a group of automorphisms of $R$ and as $K_{V}\leq\mathrm{Aut}_{\mathcal{F}^{*}_{\Lambda}(q)}^{0}(S)=\mathrm{Aut}_{\mathcal{F}}^{0}(S)$, we deduce that ${K_{V}}|_{R}\in\mathrm{Aut}_{\mathcal{F}}(R)$. It follows that $\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(R)\cong\mathrm{GL}_{2}(q)$. By similar arguments, we see that $N_{\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(R)}(\mathrm{Aut}_{S}(R))=N_{\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F}^{*}_{\Lambda}(q))}(R)}(\mathrm{Aut}_{S}(R))$. Set $B:=N_{\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(R)}(\mathrm{Aut}_{S}(R))$ and
@@
Then $Z(Y)=Z(Y^{*})=Z(B)$ has order $q-1$ and $X,X^{*}\leq C_{\mathrm{Aut}(R)}(Z(Y))=:C$. Let $t$ be the unique involution in $Z(Y^{*})$ so that $t\in Z(X^{*})$ and $R=[R,t]\times C_{R}(t)$. Then $C$ normalises $[R,t]$ and $C_{R}(t)$. Moreover, both $X$ and $X^{*}$ act faithfully on $[R,t]$ and centralise $C_{R}(t)$. An application of Lemma 5.17 reveals that $X\widehat{C}=X^{*}\widehat{C}$ where $\widehat{C}:=C_{C}([R,t])$.
Let $c$ be a $p$-element in $\widehat{C}$. Then $c$ acts trivially on $[R,t]$ and $[c,C_{R}(t)]<C_{R}(t)$. Since $Z(Y)$ acts irreducibly on $C_{R}(t)$ and commutes with $c$, we conclude that $[c,C_{R}(t)]=1$ so that $c=1$. Hence, $\widehat{C}$ is a $p^{\prime}$-group. We claim that $\widehat{C}$ is cyclic of order at most $q-1$. Note that $\widehat{C}$ acts faithfully on $C_{R}(t)$, $Z(Y)\cap\widehat{C}=\langle t\rangle$ and $\mathrm{Aut}_{Z(Y)}(C_{R}(t))$ is cyclic of order $\frac{q-1}{2}$. Moreover, $\mathrm{Aut}_{Y}(C_{R}(t))$ is cyclic of order $q-1$, so contains a Singer cycle of $\mathrm{GL}_{m}(p)$, and centralises $\mathrm{Aut}_{Z(Y)}(C_{R}(t))$. From these points we conclude, using $q>p$, that $C_{\mathrm{GL}_{m}(p)}(\mathrm{Aut}_{Z(Y)}(C_{R}(t)))=\mathrm{Aut}_{Y}(C_{R}(t))$ has order $q-1$. But $\mathrm{Aut}_{\widehat{C}}(C_{R}(t))$ centralises $\mathrm{Aut}_{Z(Y)}(C_{R}(t))$ and we therefore $C_{R}(t)$ is cyclic of order at most $q-1$. Now, $X$ acts on $\widehat{C}$ and as $X\cong\operatorname{SL}_{2}(q)$, we deduce that $X$ centralises $\widehat{C}$. But then $X=O^{p^{\prime}}(X\widehat{C})$ and so $X=X^{*}$ as desired. ∎
@@
Proofs of Theorems 1.1 and 1.2.
Suppose that $S=S_{n}(q)$ for some $11\leq n\leq p-$ 1 and $q$ is a power of $p$. Assume that $\mathcal{F}$ is a saturated fusion system on $S$ with $O_{p}(\mathcal{F})=1$. If $q=p$ then $S$ has an elementary abelian subgroup of index $p$ and so $\mathcal{F}$ is known by [Oli14, COS17]. Hence we may assume that $q>p$. If $n=1$, then Proposition 7.1 shows that $\mathcal{F}$ is realised by a group $G$ with $F^{*}(G)\cong\mathrm{PSL}_{3}(q)$. In particular, $\mathcal{F}$ is a polynomial fusion system in this case. So assume that $n\geq 2$. Then Propositions 6.1 and 6.3 gives the possibilities for the configuration of essential subgroups and their automisers. In the exceptional case that $S=S_{2}(9)$ and $O^{3^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))\cong 2^{.}\mathrm{PSL}_{3}(4)$, Proposition 7.2 shows that $\mathcal{F}$ is one of the two exotic Henke-Shpectorov systems. Thus if $V$ is $\mathcal{F}$-essential then $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))\cong\mathrm{(P)SL}_{2}(q)$ and $V\cong V_{n}(q)$. Furthermore, in this situation either $Q$ or $R$ is also $\mathcal{F}$-essential. The uniqueness of these fusion systems is then addressed in Theorem 7.7 and Theorem 7.8. The only remaining case from Proposition 6.1 is that $\mathcal{E}(\mathcal{F})=R^{S}$, and this is the subject of Theorem 7.4.

### v2

@@ 7. Uniqueness of the saturated fusion systems on $S$
Lastly suppose that $S\cong S_{\Lambda}(q)$ and $\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$. Let $k_{R}$ be the lift to $\mathrm{Aut}_{\mathcal{F}}(S)$ of an element of order q1q-1 in $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))$ so that $k_{R}$ acts on $V$. Let $k$ to be any element of order q1q-1 in $N_{\mathrm{Aut}_{\mathcal{F}}(V)}(\mathrm{Aut}_{S}(V))$ in the same Hall $p^{\prime}$-subgroup of $N_{\mathrm{Aut}_{\mathcal{F}}(V)}(\mathrm{Aut}_{S}(V))$ as $k_{R}|_{V}$, whose lift $k_{V}$ to $\mathrm{Aut}_{\mathcal{F}}(S)$ normalises $R=S_{0}Z(S)$. Then $K_{R}:=\langle k_{R}\rangle$ normalises $K_{V}:=\langle k_{V}\rangle$ and so to prove that $|\mathrm{Out}_{\mathcal{F}}^{0}(S)|=(q-1)^{2}$ as in (4), it suffices to verify $K_{R}\cap K_{V}=1$. By Proposition 6.3, $V$ is isomorphic to $\Lambda(q)$ as an $\mathbb{F}_{p}O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$-module. Let $W$ be the unique proper non-trivial $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$-submodule of $V$. Then $\mathrm{Aut}_{\mathcal{F}}(S)$ normalises $W$ since $V$ is characteristic in $S$ and $\mathrm{Aut}_{\mathcal{F}}(S)|_{V}$ normalises $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$. Let $t\in K_{R}\cap K_{V}$. Then since $t$ is the lift of an element of $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))$, $t$ acts trivially on $C_{V}(S)W/W=C_{R}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R)))W/W\cong C_{R}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R)))$. On the other hand, since $V/W$ is isomorphic as a $\mathbb{F}_{p}O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$-module to a natural $\operatorname{SL}_{2}(q)$-module, every non-trivial $p^{\prime}$-element in $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))$ acts fixed point freely on $V/W$. This implies that $t=1$ so that $K_{R}\cap K_{V}=1$, as desired.
@@
Let $\mathfrak{D}$ be a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}(S)$ containing a Hall $p^{\prime}$-subgroup $K$ of $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$. Conjugating by $\mathrm{Aut}(S)$ if necessary we may assume that $\mathrm{Aut}_{\mathcal{F}}^{0}(S)\leq\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}(S)$ and that $\mathfrak{D}$ is a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}(S)$. Recall the map $\delta:\mathrm{Aut}(S)\to\Delta=(\mathrm{Aut}(S,S/V),\mathrm{Aut}(S,Z(S)))$ from Notation 2.25. By Lemma 2.26, $\delta$ is injective upon restriction to $\mathfrak{D}$. Thus in each case, there is $k\in\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ cyclic of order q1q-1 determined by the lift of an element in $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(X))}(\mathrm{Aut}_{S}(X))$, for $X\in\{Q,R\}$ which determines a unique subgroup of $\mathfrak{D}$. Similarly if $V\in\mathcal{E}(\mathcal{F})$, then the image under $\delta$ of lift to $S$ of an element of $\frac{q-1}{(n,2)}$ in $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))}(\mathrm{Aut}_{S}(V))$ uniquely determines the actions on $S/V$ and $Z(S)$. We conclude that $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ is uniquely determined. ∎
@@
If $S=S_{n}(q)$ for some $2\leqslant n\leqslant p-1$ and $\mathcal{E}(\mathcal{F})=R^{S}$ then $\mathcal{F}$ is isomorphic to a $p^{\prime}$-index subsystem of $\mathcal{F}^{*}(n,q,R)_{P}$.
@@
In all remaining cases, we may assume that $V\in\mathcal{E}(\mathcal{F})$ by Propositions 6.1 and 6.2. Recall the group $D^{\dagger}$ from Notation 2.1.
@@
[…]
@@
If $S=S_{n}(q)$ and $\mathcal{E}(\mathcal{F})=\{V\}\cup Q^{S}$ then $O^{p^{\prime}}(\mathcal{F})\cong O^{p^{\prime}}(\mathcal{F}^{*}(n,q,Q))$ and $\mathcal{F}$ is isomorphic to a $p^{\prime}$-index subsystem of $\mathcal{F}^{*}(n,q,Q)$.
@@
Write $X:=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(Q))$ and $X^{*}:=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,Q)}(Q))$ so that $N_{X}(\mathrm{Aut}_{S}(Q))=N_{X^{*}}(\mathrm{Aut}_{S}(Q))$. By Proposition 2.9 and Corollary 2.13, $\mathrm{Aut}(Q)=CH$ where $H\cong\Gamma\mathrm{L}_{2}(q)$ and $C$ is a normal $p$-subgroup of central automorphisms of $Q$. It follows immediately that $XC=X^{*}C$. Let $\tau$ be an involution in $N_{X}(\mathrm{Aut}_{S}(Q))$, so $X=\mathrm{Inn}(Q)C_{X}(\tau)$ and $X^{*}=\mathrm{Inn}(Q)C_{X^{*}}(\tau)$. Since $C_{C}(\tau),Q,\tau]=1$, the Three Subgroups Lemma implies that $Q,C_{C}(\tau)]=[[\tau,Q]Z(Q),C_{C}(\tau)]=1$, and so $C_{C}(\tau)=1$. We deduce that $C_{X}(\tau)=C_{XC}(\tau)=C_{X^{*}C}(\tau)=C_{X^{*}}(\tau)$, so $X=X^{*}$ and $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(Q))=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,Q)}(Q))$, as desired. ∎
@@
If $S=S_{n}(q)$ and $\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$ then $O^{p^{\prime}}(\mathcal{F})\cong O^{p^{\prime}}(\mathcal{F}^{*}(n,q,R))$ and $\mathcal{F}$ is isomorphic to a $p^{\prime}$-index subsystem of $\mathcal{F}^{*}(n,q,R)$.
@@
By Propositions 7.5 and 7.6 it suffices to show that $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ uniquely determines $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))$. Set $X:=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))$ and $X^{*}=O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}(R))$ and observe that $N_{X}(\mathrm{Aut}_{S}(R))$ is contained in the restriction of $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ to $R$. Upon demonstrating that $N_{X}(\mathrm{Aut}_{S}(R))=N_{X^{*}}(\mathrm{Aut}_{S}(R))$, the result will follow from Lemma 5.17. Recall that $\mathrm{Aut}_{\mathcal{F}}^{0}(S)=\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}^{0}(S)$ and let $k$ be the lift of an element of order q1q-1 in $N_{X}(\mathrm{Aut}_{S}(R))$ to $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$. Then $k$ acts as $\mu^{-1}$ on $S/V$ and as $\mu$ on $C_{V}(S)$. By Lemma 2.26, $\delta$ is injective upon restricting to a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{\mathcal{F}^{*}(n,q,R)}^{0}(S)$, so the restriction of $k$ to $R$ lies in $N_{X^{*}}(\mathrm{Aut}_{S}(R))$. But then $N_{X}(\mathrm{Aut}_{S}(R))=\langle\mathrm{Aut}_{S}(R),k\rangle=N_{X^{*}}(\mathrm{Aut}_{S}(R))$ and so Lemma 5.17 applies. This completes the proof. ∎
@@
If $S=S_{\Lambda}(q)$ and $\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$ then $O^{p^{\prime}}(\mathcal{F})\cong O^{p^{\prime}}(\mathcal{F}^{*}_{\Lambda}(q))$ and $\mathcal{F}$ is isomorphic to a $p^{\prime}$-index subsystem of $\mathcal{F}^{*}_{\Lambda}(q)$.
@@
By Propositions 7.5 and 7.6 it suffices to show that $\mathrm{Aut}_{\mathcal{F}}^{0}(S)$ uniquely determines $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))$. Let $k$ be the lift to $\mathrm{Aut}_{\mathcal{F}}(S)$ of an element of $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(R))}(\mathrm{Aut}_{S}(R))$ of order q1q-1. Since $\mathrm{Aut}_{\mathcal{F}}^{0}(S)=\mathrm{Aut}_{\mathcal{F}^{*}_{\Lambda}(q)}^{0}(S)$, $k$ lies in a Hall $p^{\prime}$-subgroup of $N_{\mathrm{Aut}_{\mathcal{F}^{*}_{\Lambda}(q)}^{0}(S)}(R)$ and there is a subgroup $K_{V}$ of order q1q-1 satisfying
@@
Thus $K_{V}$ restricts faithfully to a group of automorphisms of $R$ and as $K_{V}\leq\mathrm{Aut}_{\mathcal{F}^{*}_{\Lambda}(q)}^{0}(S)=\mathrm{Aut}_{\mathcal{F}}^{0}(S)$, we deduce that ${K_{V}}|_{R}\in\mathrm{Aut}_{\mathcal{F}}(R)$. It follows that $\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(R)\cong\mathrm{GL}_{2}(q)$. By similar arguments, we see that $N_{\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(R)}(\mathrm{Aut}_{S}(R))=N_{\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F}^{*}_{\Lambda}(q))}(R)}(\mathrm{Aut}_{S}(R))$. Set $\mathbb{B}:=N_{\mathrm{Aut}_{O^{p^{\prime}}(\mathcal{F})}(R)}(\mathrm{Aut}_{S}(R))$ and
@@
Then $Z(Y)=Z(Y^{*})=Z(\mathbb{B})$ has order q1q-1 and $X,X^{*}\leq C_{\mathrm{Aut}(R)}(Z(Y))=:C$. Let $t$ be the unique involution in $Z(Y^{*})$ so that $t\in Z(X^{*})$ and $R=[R,t]\times C_{R}(t)$. Then $C$ normalises $[R,t]$ and $C_{R}(t)$. Moreover, both $X$ and $X^{*}$ act faithfully on $[R,t]$ and centralise $C_{R}(t)$. An application of Lemma 5.17 reveals that $X\widehat{C}=X^{*}\widehat{C}$ where $\widehat{C}:=C_{C}([R,t])$.
[…]
@@
Proofs of Theorems 1.1 and 1.3.
Suppose that $S=S_{n}(q)$ for some $1\leq n\leq p-1$ and $q$ is a power of $p$. Assume that $\mathcal{F}$ is a saturated fusion system on $S$ with $O_{p}(\mathcal{F})=1$. If $q=p$ then $S$ has an elementary abelian subgroup of index $p$ and so $\mathcal{F}$ is known by [OLI14, COS17]. Hence we may assume that $q>p$. If $n=1$, then Proposition 7.1 shows that $\mathcal{F}$ is realised by a group $G$ with $F^{*}(G)\cong\mathrm{PSL}_{3}(q)$. In particular, $\mathcal{F}$ is a polynomial fusion system in this case. If $p=2$ then $S_{\Lambda}(q)$ has nilpotency class two and Theorem 1.3(1) follows from [VAN24b, Proposition 4.2]
So assume that $p$ is odd and $n\geqslant 2$. Then Propositions 6.1 and 6.3 give the possibilities for the configuration of essential subgroups and their automisers. In the exceptional case that $S=S_{2}(9)$ and $O^{3^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))\cong 2^{.}\mathrm{PSL}_{3}(4)$, Proposition 7.2 shows that $\mathcal{F}$ is one of the two exotic Henke–Shpectorov systems. Thus if $V$ is $\mathcal{F}$-essential then $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(V))\cong\mathrm{(P)SL}_{2}(q)$ and $V\cong V_{n}(q)$. Furthermore, in this situation either $Q$ or $R$ is also $\mathcal{F}$-essential. The uniqueness of these fusion systems is then addressed in Theorem 7.7 and Theorem 7.8. The only remaining case from Proposition 6.1 is that $\mathcal{E}(\mathcal{F})=R^{S}$, and this is the subject of Theorem 7.4.